If the future value of an ordinary, 8-year annuity is $6,400 and interest rates are 8.0 percent, what is the future value of the same annuity due?
1
Expert's answer
2018-07-06T09:07:08-0400
If the future value of an ordinary, 8-year annuity is $6,400 and interest rates are 8.0 percent, then:
6,400 = P*((1 + 0.08)^8 - 1)/0.08,
P = 6,400*0.08/(1.08^8 - 1) = $601.7.
So, the future value of the same annuity due is:
FV = (1 + 0.08)*601.7*((1 + 0.08)^8 - 1)/0.08 = $6,912.06.